# Taking the Next Step with Generative Artificial Intelligence: The Transformative Role of Multimodal Large Language Models in Science Education

### v1

@@ Abstract.
The integration of Artificial Intelligence (AI), particularly Large Language Model (LLM)-based systems, in education has shown promise in enhancing teaching and learning experiences. However, the advent of Multimodal Large Language Models (MLLMs) like GPT-4 with vision (GPT-4V), capable of processing multimodal data including text, sound, and visual inputs, opens a new era of enriched, personalized, and interactive learning landscapes in education. Grounded in theory of multimedia learning, this paper explores the transformative role of MLLMs in central aspects of science education by presenting exemplary innovative learning scenarios. Possible applications for MLLMs could range from content creation to tailored support for learning, fostering competencies in scientific practices, and providing assessment and feedback. These scenarios are not limited to text-based and uni-modal formats but can be multimodal, increasing thus personalization, accessibility, and potential learning effectiveness. Besides many opportunities, challenges such as data protection and ethical considerations become more salient, calling for robust frameworks to ensure responsible integration. This paper underscores the necessity for a balanced approach in implementing MLLMs, where the technology complements rather than supplants the educator’s role, ensuring thus an effective and ethical use of AI in science education. It calls for further research to explore the nuanced implications of MLLMs on the evolving role of educators and to extend the discourse beyond science education to other disciplines. Through the exploration of potentials, challenges, and future implications, we aim to contribute to a preliminary understanding of the transformative trajectory of MLLMs in science education and beyond.
@@ 1. Introduction
Science education encompasses a broad spectrum of activities, from acquiring scientific knowledge and engaging in scientific practices to effective communication about scientific findings and ideas (US: NRC, (2012)). These competencies are designed to prepare students for the complex, multifaceted challenges they will encounter in today’s world (OECD,, 2018) and, therefore, are often part of 21st-century skills frameworks (e.g., Kay and Dardis, (2016); for science education: McComas, (2014)). The essence of science learning is inherently multimodal. It necessitates that students engage in a range of activities within different modalities: reading and writing scientific texts, interpreting diagrams, drawing models, flowcharts, and diagrams, analyzing and visualizing data, designing solutions, synthesizing information across various modalities, and communicating scientific ideas. These activities shape the understanding of scientific knowledge and the development of appropriate domain-specific competencies (Lemke,, 1998; Van Leeuwen and Kress,, 1990; Doran et al.,, 2021). Additionally, it requires students to seamlessly shift between these different modes. Representations such as images, graphics, tables, technical symbols, formulas, equations, process workflows, graphs, and video as well as audio data, are used in the learning process. This multimodal nature of science learning demands science educators to facilitate learning with multimodal materials and provide students with opportunities to engage in multimodal activities.
@@
Beyond the multimodal nature of science itself, the general effectiveness of multimodal learning can serve as a second argument for fostering multimodality in science education. ”Multimodal learning” (as defined and used herein) has a cognitive foundation based on the multimedia learning theory (CTML) posited by Mayer (e.g., Mayer, (1997, 2021)). Combining multimodal representations like text and images can enhance knowledge acquisition, improving hence the integration into a coherent multi-faceted mental model. Schnotz and Bannert, (2003) have provided strong evidence for multimodal approaches in (science) education and state, in accordance with (Mayer,, 1997), that learning with multimodal material is highly effective. The real-time generation of multimodal material offers many opportunities for adaptive learning. Therefore, multimodality is not only an integral part of science education itself but, from the perspective of cognitive science, can be harnessed to build effective learning environments.
@@
With the latest advances in Artificial Intelligence (AI), specifically in generative AI, Large Language Models (LLMs) have become increasingly popular in the general educational context (cf. Seßler et al., (2023); Moore et al., (2022); Abdelghani et al., 2023a ) and science education (cf. Bewersdorff et al., 2023a ; Küchemann et al., (2023)), with strong potential to advance multimodal learning. The use of LLMs in education has grown mainly due to their ability to reason about, generate, and manipulate text data. Since LLMs merely focus on the textual mode, they generally do not capture the variety of multimodal representations in science. With the emergence of Multimodal Large Language Models (MLLMs), we are now able to meet the necessity for multimodality in science education. Educators are able to use MLLM, such as GPT-4V, to read scoring rubrics and grade students’ drawn models (Lee and Zhai,, 2023). MLLMs can process and create content not only in text but across multiple forms of data, including images, audio, and video (Bommasani et al.,, 2021). They can also transform unstructured data into structured content and vice versa (Borisov et al., 2022b, ). This shift from text-centric approaches towards multimodal content generation opens a new era in science education, where the capabilities of LLMs are expanded to meet the multimodal demands of science education.
@@
In this article, we articulate the potential of MLLMs, specifically in the field of science education. Building upon prior research on LLMs in education (e.g., Kasneci et al., (2023); Zhai, (2023); Lee et al., 2023b ) we assess the potential of MLLMs in generally reshaping educational processes. Guided by Mayer’s CTML, we explore the potential of MLLMs in exemplary scenarios of multimodality in instructional strategies and design, student engagement, assessment, and feedback. We will also discuss the challenges related to implementing and applying MLLMs in the classroom, such as data protection and ethical concerns, which become even more pressing with multimodal systems. These multimodal systems are potentially capable of processing not only (mostly) anonymous textual data but also, depending on the specific use cases, personal data like voice recordings or images of handwriting. Tackling these issues is crucial to responsibly integrating MLLMs in learning environments and fully realizing their potential to improve education (Lee et al., 2023b, ).
@@ 2.1. Core Elements of Science Education
Science education is crucial in preparing informed citizens capable of navigating an increasingly complex world (OECD,, 2018). This education involves imparting a comprehensive understanding of core scientific concepts and crosscutting ideas essential for building a solid foundation in science (NRC,, 2012). Alongside this robust content knowledge, the development of scientific thinking and practical skills is critical. Engaging students in scientific inquiry and hands-on investigations not only cultivates their critical thinking but also equips them with practical abilities essential for scientific practices (Bybee,, 1997; Flick,, 1993). Moreover, such an approach may also foster their creativity, enabling them to devise effective solutions for ill-structured problems (Hathcock et al.,, 2015). Central to science education is also the emphasis on effective communication skills. Students should be able to articulate and communicate complex scientific ideas with clarity and precision (NRC,, 2012). This significance was particularly highlighted during the recent COVID-19 pandemic, where the adept communication of scientific results played a crucial role in informing the public and guiding decision-making processes in policy-making institutions (Besançon et al.,, 2021).
@@
Central tasks for educators are creating engaging content and empowering students’ learning. In science education, the latter involves not only fostering the acquisition of content knowledge but also encouraging engagement in scientific practices and the communication of scientific findings, insights, and ideas. Assessment and providing effective feedback is a third crucial task of educators. These three central aspects of science education will guide the presentation of the exemplary potentials of MLLMs in chapter 3.
2.2. Large Generative AI Models
@@
Large Language Models. LLMs are a subset of AI models focusing on processing and generating human-like text, enabling thus a wide range of applications from content creation to problem-solving (Brown et al.,, 2020). Those deep learning architectures are characterized by their ability to understand context, generate coherent responses, and provide textual content in real-time.
@@
Incorporating human feedback through the training of reward models and applying reinforcement learning in the fine-tuning step (Ouyang et al.,, 2022; Lee et al., 2023c, ) has significantly enhanced the generation of human-like language that is aligned with social norms and goals and has led to the success of today’s popular language models by OpenAI, ChatGPT and GPT-4 (OpenAI,, 2023). The PaLM models (Chowdhery et al.,, 2022; Anil et al.,, 2023) from Google are setting a focus on a smaller and denser architecture, achieving similar performance. Also, in the open-source landscape, there are strong competitors like Meta’s LLaMA model (Touvron et al., 2023a, ; Touvron et al., 2023b, ), and the Falcon-40B architecture (Penedo et al.,, 2023). These models have sparked further development, with researchers fine-tuning them and creating open-source variants like Vicuna (Zheng et al.,, 2023) and Alpaca (Taori et al.,, 2023).
@@
The impact of these LLMs has been profound and led to novel advances across various industries. They have enabled innovative approaches in areas such as healthcare (Chintagunta et al.,, 2021; Enarvi et al.,, 2020), finance (Dowling and Lucey,, 2023), journalism (Pavlik,, 2023), creative writing (Yuan et al.,, 2022), and even in scientific discovery (Romera-Paredes et al.,, 2023), demonstrating the strength and versatility of LLMs. The advances in the generation of human-like texts have opened up new opportunities and challenges in different domains (Kaddour et al.,, 2023). This includes a meaningful and constructive application in educational contexts, considering remaining issues like correct mathematical reasoning (Imani et al., 2023a, ), or possible hallucinations in the output (Ji et al., 2023a, ). Research continues to address these challenges while also improving their explainability to increase trust and reliability in the model outputs (Wu et al., 2023c, ). While LLMs have significantly impacted text-based learning (e.g., Bewersdorff et al., 2023a ; Seßler et al., (2023); Küchemann et al., (2023)), the advent of MLLMs promises to extend these benefits to include visual, auditory, and other sensory data modalities.
@@
Multimodal Large Language Models. Building upon the foundation of LLMs, MLLMs are a further advancement considering their ability to process and generate content across various data types, including text, images, audio, and video (Bommasani et al.,, 2021). However, these models are built to understand, interpret, and respond to multiple modalities, enriching context information beyond merely text-based capabilities. Much initial research focuses on the combination of visual and textual understanding as a multimodal input, yet they predominantly generate output in text form only (Alayrac et al.,, 2022; Zhu et al.,, 2023; Huang et al.,, 2023; Ye et al.,, 2023). To accomplish this joint understanding of textual and visual input, developers often integrate existing image encoders with pre-trained LLMs. They enable this combination by incorporating additional modules that project the different data forms into a shared embedding space (Li et al., 2023b, ; Liu et al.,, 2023; Zhang et al., 2023c, ).
@@
However, the capabilities of MLLMs are not limited to images and text; they can also process combinations of text and video data (Li et al., 2023b, ; Maaz et al.,, 2023). For example, this can be achieved by dissecting video into audio and a series of images, then encoding and embedding each element separately to capture the context of the video format (Zhang et al., 2023c, ). Taking a more comprehensive approach, multimodal inputs can also move beyond mere dual-modality and include additional sensor data (Driess et al.,, 2023) or speech (Chen et al.,, 2023). An example of this advancement is PandasGPT (Su et al.,, 2023), which integrates six different modalities as input for LLMs.
@@
Enhancements in multimodal capabilities, such as more advanced structures for audio processing (Zhang et al., 2023a, ), have the potential to significantly boost performance. The concept of a singular model capable of handling text, images, audio, and video for both input and output is still relatively nascent. Recently, Wu et al. published NExT-GPT as a general-purpose any-to-any open-source MLLM (Wu et al., 2023b, ) and OpenAI released GPT-4 Vision (OpenAI,, 2023) and GPT-4 Turbo [^1] https://chatopenai.de/gpt-4-turbo/, marking an important further development in this research. Also, Google announced an new LLM-based AI system they named Gemini, which is, according to Google ”built from the ground up for multimodality - reasoning seamlessly across text, images, video, audio, and code” (DeepMind,, 2023). This kind of reasoning in combination with other types of deep learning and explainability approaches (Wu et al.,, 2020; Borisov et al., 2022a, ; Rombach et al.,, 2022) has the potential to advance science (Wong et al.,, 2023) and education. Often, these AI systems are labeled Multimodal Foundation Models (Li et al., 2023a, ). Following Fu et al., (2023), we will use the term ’Multimodal Large Language Models’ in our paper.
2.3. Adaptive Multimodal Learning
@@
Human cognitive processing of information occurs through various channels, each handling different forms of information, such as spoken or written language or images. The insight that combining multiple modalities can enhance knowledge acquisition is documented in Mayer’s theory of multimedia learning (Mayer,, 1997, 2021) as well as in Schnotz & Bannert’s theory of text-picture integration (Schnotz and Bannert,, 2003). Based on cognitive models of information processing, such as Paivio’s Dual Coding Theory (Paivio,, 1991), the so-called multimedia effect is the well-supported hypothesis that active processing of both textual and pictorial information enhances integration into a coherent mental model (for a recent meta-analysis of the multimedia effect, see Hu et al., (2021)). Knowledge acquisition becomes more complete and less ambiguous and, thus, more effective when learners construct mental models from different modalities. Different representations can complement each other or even lead to a better overall understanding. For example, an image of a human skeleton or the depiction of an enzymatic reaction could be more effective and easier to understand than a written description, particularly when considering the time required to process the information. However, the verbal representation of knowledge can complement the image, especially when terms presented in text are unknown. Statements and deeper thoughts can also be represented visually, although not unambiguously.
@@
According to the Cognitive Theory of Multimedia Learning (CTML), multimodal representation enables the selection, organization, and integration of learning content into a coherent mental model (Mayer,, 2021). Learners first select relevant information from the presented material to focus their attention on. The selected information is then organized in the learner’s mind into coherent structures, and finally, learners integrate the organized information with their existing knowledge to make sense of it. For instance, according to the signaling principle (cf. van Gog, (2014)), learners often have difficulty selecting important information in learning materials, so they need to be supported in their search to avoid unnecessary cognitive load, especially in text-image relations (for a meta-analysis of different studies on the signaling principle, see Richter et al., (2016)). Previous research has shown that signaling, i.e., separate selection aids, can hinder learning, especially for learners with at least sufficient prior knowledge (e.g., Richter and Scheiter, (2019)).
@@
Second, learners have to organize their learning materials, i.e., text and images, according to their needs to keep their cognitive demands as low as necessary for efficient learning processes to take place (Mayer and Moreno,, 2003). Third, according to Mayer, (2021), the integration of multimodal representations requires generative activities on the part of the learner. It is, therefore, essential that multimodal representations are actively transformed into coherent mental models by the student.
As will be discussed below in relation to specific science education contexts, MLLMs offer a potential added value to this three-step process by enabling learners and educators to adapt multimodal representations themselves. As the CTML refers to the processing of multimodal information in the context of teaching-learning processes (cf. Mayer, (2021)), it provides a suitable conceptual framework for the following discussion of the potential of MLLMs in (science) education.

### v2

@@ Abstract.
The integration of Artificial Intelligence (AI), particularly Large Language Model (LLM)-based systems, in education has shown promise in enhancing teaching and learning experiences. However, the advent of Multimodal Large Language Models (MLLMs) like GPT-4 with vision (GPT-4V), capable of processing multimodal data including text, sound, and visual inputs, opens a new era of enriched, personalized, and interactive learning landscapes in education. Grounded in theory of multimedia learning, this paper explores the transformative role of MLLMs in central aspects of science education by presenting exemplarily powerful learning scenarios. Possible applications for MLLMs could range from content creation to tailored support for content learning, fostering competencies in scientific practices, and providing assessment and feedback. These scenarios are not limited to text-based and uni-modal formats but can be multimodal, increasing personalization, accessibility, and potential learning effectiveness. Besides its potential, challenges such as data protection and ethical considerations become more salient, calling for robust frameworks to ensure responsible integration. This paper underscores the necessity for a balanced approach in implementing MLLMs, where the technology complements rather than supplants the teacher’s role, ensuring an effective and ethical use of AI in science education. It calls for further research to explore the nuanced implications of MLLMs on the evolving role of teachers and to extend the discourse beyond science education to other disciplines. Through the exploration of potentials, challenges, and future implications, we aim to contribute to a preliminary understanding of the transformative trajectory of MLLMs in science education and beyond.
@@ 1. Introduction
Science education encompasses a broad spectrum of activities, from acquiring scientific knowledge and engaging in scientific practices to effective communication about scientific findings and ideas (Council et al.,, 2012). These aspects are designed to prepare students for the complex, multifaceted challenges they will encounter in today’s world (OECD,, 2018) and, therefore, are often part of 21st-century skills frameworks (e.g., science education: (McComas,, 2014)). The essence of science learning is inherently multimodal. It necessitates that students engage in a range of activities: reading and writing scientific texts, interpreting diagrams, drawing models, analyzing data, designing solutions, and synthesizing information across various modalities. These activities shape the understanding of scientific knowledge and the development of appropriate domain-specific skills (Lemke,, 1998; Van Leeuwen and Kress,, 1990; Doran et al.,, 2021). Additionally, it requires students to seamlessly shift between these different modes. Representations such as images, graphics, tables, technical symbols, formulas, equations, graphs, and video as well as audio data, are used in the learning process. This multimodality nature of science learning demands science educators to facilitate learning with multimodal materials and provide students opportunities to engage in multimodal activities.
@@
”Multimodal learning” (as defined and used herein) has a cognitive foundation based on the multimedia learning theory (CTML) posited by Mayer (e.g., Mayer, 1997, 2021). Combining multimodal representations like text and images can enhance knowledge acquisition, improving the integration into a coherent mental model. Mayer, (1997); Schnotz and Bannert, (2003) have provided strong evidence for multimodal approaches in science education and state that, in accordance with Mayer, (1997) that learning with multimodal material is highly effective. The real-time generation of multimodal material offers many opportunities for adaptive learning. Therefore, multimodality is not only an integral part of science education itself but, from the perspective of cognitive science, can be harnessed to build effective learning environments.
@@
With the latest advances in Artificial Intelligence (AI), Large Language Models (LLMs) have become increasingly popular in the general educational context cf .(Seßler et al.,, 2023; Moore et al.,, 2022; Abdelghani et al., 2023a, ) and science education cf. (Bewersdorff et al.,, 2023; Küchemann et al.,, 2023), with strong potential to advance multimodal learning. The use of LLMs in education has grown mainly due to their ability to understand and generate text data (Lee et al., 2023a, ). Since LLMs merely focus on the textual mode, they generally do not capture the variety of multimodal representations in science. With the emergence of MLLMs, we are now able to meet the necessity for multimodality in science education. Researchers are able to use MLLM, such as GPT-4V, to read scoring rubrics and grade students’ drawn models (Lee and Zhai,, 2023). MLLMs can process and create content not only in text but across multiple forms of data, including images, audio, and video (Bommasani et al.,, 2021). This shift from text-centric approaches opens a new era in science education, where the capabilities of LLMs are expanded to meet the multimodal demands of science education.
@@
In this article, we articulate the potential of MLLMs, specifically in the field of science education. Building upon prior research on LLMs in education (e.g., (Kasneci et al.,, 2023; Zhai,, 2023; Lee et al., 2023b, ) we assess the potential of MLLMs in generally reshaping educational processes. Guided by Mayer’s CTML, we explore the potential of MLLMs in exemplary scenarios of multimodality in instructional strategies and design, enhanced student engagement, assessment, and feedback. We will also discuss the challenges related to implementing and applying MLLMs in the classroom, such as data protection and ethical concerns, which become even more pressing with multimodal systems. These multimodal systems are potentially capable of processing not only (mostly) anonymous textual data but also, depending on the specific use cases, personal data like voice recordings or handwriting images. Tackling these issues is crucial to responsibly integrating MLLMs and fully realizing their potential to improve education (Lee et al., 2023b, ).
@@ 2.1. Core Elements of Science Education
Science education is crucial in preparing informed citizens capable of navigating an increasingly complex world (OECD,, 2018). This preparation involves imparting a comprehensive understanding of core scientific concepts and crosscutting ideas essential for building a solid foundation in science (Council et al.,, 2012). Alongside this robust content knowledge, the development of scientific thinking and practical skills is critical. Engaging students in scientific inquiry and hands-on investigations not only cultivates their critical thinking but also equips them with practical abilities essential for scientific practices (Bybee,, 1997; Flick,, 1993). Central to science education is also the emphasis on effective communication skills. Students should be able to articulate complex scientific ideas with clarity and precision.
@@
Central tasks for teachers are the creation of engaging content and empowering students’ learning. Empowering students’ learning in science education involves not only fostering the acquisition of content knowledge but also encouraging engagement in scientific practices and the communication of scientific findings and ideas. Assessment and providing effective feedback is a third crucial task of teachers. These central aspects of science education will guide the presentation of the exemplary potentials of MLLMs in chapter 3.
2.2. Large Language Models
@@
Large Language Models (LLMs) are a subset of Artificial Intelligence (AI) focusing on processing and generating human-like text, enabling a wide range of applications from content creation to problem-solving (Brown et al.,, 2020). Those deep learning architectures are characterized by their ability to understand context, generate coherent responses, and provide textual content in real time.
@@
Incorporating human feedback through the training of reward models and applying reinforcement learning in the fine-tuning step (Ouyang et al.,, 2022) has significantly enhanced human-like language generation and led to the success of today’s popular language models by OpenAI, ChatGPT and GPT-4 (OpenAI,, 2023). The PaLM models (Chowdhery et al.,, 2022; Anil et al.,, 2023) from Google are setting a focus on a smaller and denser architecture, achieving similar performance. Also, in the open-source landscape, there are strong competitors like Meta’s LLaMA model (Touvron et al., 2023a, ; Touvron et al., 2023b, ), and the Falcon-40B architecture (Penedo et al.,, 2023). These models have sparked further development, with researchers fine-tuning them and creating open-source variants like Vicuna (Zheng et al.,, 2023) and Alpaca (Taori et al.,, 2023).
@@
The impact of these LLMs has been profound and led to novel advances across various industries. They have enabled innovative approaches in areas such as healthcare (Chintagunta et al.,, 2021; Enarvi et al.,, 2020), finance (Dowling and Lucey,, 2023), journalism (Pavlik,, 2023), and creative writing (Yuan et al.,, 2022), demonstrating the strength and versatility of LLMs. The advances in the generation of human-like texts have opened up new opportunities and challenges have in different domains (Kaddour et al.,, 2023). This includes a meaningful and constructive application in educational contexts, considering remaining issues like correct mathematical reasoning (Imani et al., 2023a, ), reliable student error detection (Bewersdorff et al.,, 2023), or possible hallucinations in the output (Ji et al.,, 2023). Research continues to address these challenges while also improving their explainability to increase trust and reliability in the model outputs (Wu et al., 2023c, ). While LLMs have significantly impacted text-based learning (e.g. (Bewersdorff et al.,, 2023; Seßler et al.,, 2023; Küchemann et al.,, 2023)), the advent of MLLMs promises to extend these benefits to include visual, auditory, and other sensory data modalities.
2.3. Multimodal Large Language Models
@@
Building upon the foundation of LLMs, MLLMs are a further advancement considering their ability to process and generate content across various data types, including text, images, audio, and video (Bommasani et al.,, 2021). These models are built to understand, interpret, and respond to multiple modalities, enriching context information beyond merely text-based capabilities. Much initial research focuses on the combination of visual and textual understanding as a multimodal input, yet they predominantly generate output in text form only (Alayrac et al.,, 2022; Zhu et al.,, 2023; Huang et al.,, 2023; Ye et al.,, 2023). To accomplish this joint understanding of textual and visual input, developers often integrate existing image encoders with pre-trained LLMs. They enable this combination by incorporating additional modules that project the different data forms into a shared embedding space (Li et al.,, 2023; Liu et al.,, 2023; Zhang et al., 2023b, ).
@@
However, the capabilities of MLLMs are not limited to images and text; they can also process combinations of text and video data (Li et al.,, 2023; Maaz et al.,, 2023). For example, this can be achieved by dissecting video into audio and a series of images, then encoding and embedding each element separately to capture the context of the video format (Zhang et al., 2023b, ). Taking a more comprehensive approach, multimodal inputs can also move beyond mere dual-modality and include additional sensor data (Driess et al.,, 2023) or speech (Chen et al.,, 2023). An example of this advancement is PandasGPT (Su et al.,, 2023), which integrates six different modalities as input for LLMs.
@@
Enhancements in multimodal capabilities, such as more advanced structures for audio processing (Zhang et al., 2023a, ), have the potential to significantly boost performance. The concept of a singular model capable of handling text, images, audio, and video for both input and output is still relatively nascent. Recently, Wu et al. published NExT-GPT as a general-purpose any-to-any open-source MLLM (Wu et al., 2023b, ) and OpenAI released GPT-4 Vision (OpenAI,, 2023), marking an important further development in this research. Also, Google announced an new LLM-based AI system they named Gemini, which is, according to Google ”built from the ground up for multimodality - reasoning seamlessly across text, images, video, audio, and code” (DeepMind,, 2023).
2.4. Adaptive, Multimodal Learning
@@
Human cognitive processing of information occurs through various channels, each handling different forms such as spoken or written text and images. The notion that combining multiple modalities can enhance knowledge acquisition is documented in Mayer’s theory of multimedia learning (1997, 2021) as well as in Schnotz & Bannert’s theory of text-picture integration (Schnotz and Bannert,, 2003). Based on cognitive models of information processing, such as Paivio’s Dual Coding Theory (1991) (Paivio,, 1991), the so-called multimedia effect is the well-supported hypothesis that active processing of both textual and pictorial information enhances integration into a coherent mental model (for a recent meta-analysis of the multimedia effect, see (Hu et al.,, 2021)). Knowledge acquisition becomes more complete and, thus, more effective when learners construct mental models from different modalities. Different representations can complement each other or even lead to a better overall understanding. For example, an image of a human skeleton or the depiction of an enzymatic reaction could be more effective and easier to understand than a written description, particularly considering the time required to process the information. However, the verbal representation of knowledge can complement the image, especially when the terms presented in text are unknown. Statements and deeper thoughts can also be represented visually, although not unambiguously.
@@
According to the theory of multimodal learning (CTML), multimodal representation enables the selection, organization, and integration of learning content into a coherent mental model (Mayer,, 2021). Learners first select relevant information from the presented material, both verbal and visual, to focus their attention on. The selected information is then organized in the learner’s mind into coherent structures, and finally, learners integrate the organized information with their existing knowledge to make sense of it. For instance, according to the signaling principle cf. (van Gog,, 2014), learners often have difficulty selecting important information in learning materials, so they need to be supported in their search to avoid unnecessary cognitive load, especially in text-image relations (for a meta-analysis of different studies on the signaling principle, see (Richter et al.,, 2016)). Previous research has shown that signaling, i.e., separate selection aids, can hinder learning, especially for learners with at least sufficient prior knowledge (e.g., (Richter and Scheiter,, 2019)).
@@
Second, learners have to organize their learning materials, i.e., text and images, according to their needs to keep their cognitive demands as low as necessary for efficient learning processes to take place (Mayer and Moreno,, 2003). Third, according to (Mayer,, 2021), the integration of multimodal representations requires generative activities on the part of the learner. It is therefore essential that multimodal representations are actively transformed into coherent mental models. As will be discussed below in relation to specific science education contexts, MLLMs offer a potential added value to this three-step process by enabling learners and teachers to adapt multimodal representations themselves.
As the Cognitive Theory of Multimedia Learning (CTML) refers to the processing of multimodal information in the context of teaching-learning processes cf. (Mayer,, 2021) it provides a suitable conceptual framework for discussing the potentials of MLLMs in (science) education.